OpenStudy (anonymous):

help with understanding steps in integrating (9x^2 - 12x +4)^-1

OpenStudy (anonymous):

I know the answer is 1/(6 - 9x) but the question I have has to do with: For the integrand 1/(3 x-2)^2, substitute u = 3 x-2 and du = 3 dx: = 1/3 integral 1/u^2 du

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

Yep, please

OpenStudy (anonymous):

specifically: why du = 3 dx?

OpenStudy (anonymous):

Let u = 3x -2 --> du = 3 dx This step means you take derivative of u with respect to x, the left hand side is \(\dfrac{du}{dx}\)

OpenStudy (anonymous):

the right hand side is 3.

OpenStudy (anonymous):

so that du = 3 dx

OpenStudy (anonymous):

Alright, I get that thanks! But then shouldn't the integral be S 3/u^2 rather than 1/3 S 1/u^2?

OpenStudy (anonymous):

hey,\(\dfrac {du}{dx} = 3 \rightarrow \dfrac{du}{3}= dx\) , right? in original problem, you have dx only, not 3 dx, so that you must get dx only to replace to it, right?

OpenStudy (anonymous):

Ahhh that makes a ton of sense! Thanks a lot!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Do you have to use that reasoning everytime you use u substitution?

OpenStudy (anonymous):

That is logic! not reasoning. When making some argument, you must follow the logic to know where you get the expression.

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

:)